use lines with the equation 3x+4y=12 and 6x+ry=12
a)find r if the lines are parallel.
b)find r if the lines are perpendicular

Answer by Alan3354(69443)

use lines with the equation 3x+4y=12 and 6x+ry=12
a)find r if the lines are parallel.
b)find r if the lines are perpendicular
-------------
The slope of 3x + 4y = ... is -3/4. (The right side has no effect on the slope)
--------
The slope of 6x + ry = ... is -6/r
---------------
a) parallel means the same slope
-6/r = -3/4
r = 8
----------
b) perpendicular means the slope is the negative inverse, = 4/3
-6/r = 4/3
r = -4.5